OpenStudy (anonymous):

(√(2x+3)-√(3))÷x simplify

OpenStudy (anonymous):

I believe that is the the equation simplified, unless you want it to look like these: sqrt(2 x+3)/x-sqrt(3)/x -(sqrt(3)-sqrt(2 x+3))/x http://www.wolframalpha.com/input/?i=%28 √%282x%2B3%29-√%283%29%29÷x+

OpenStudy (anonymous):

i'm pretty sure the problem is asking me to combine the two square roots. is there any way to do that?

OpenStudy (anonymous):

No, sorry!?!

OpenStudy (anonymous):

oh well, thanks anyway

OpenStudy (anonymous):

Did you see the first one?

OpenStudy (anonymous):

sqrt(2 x+3)/x-sqrt(3)/x \[\frac{ \sqrt{2x+3} }{ x } - \frac{ \sqrt{3} }{ x }\]